A random sample of 160 observations results in 104 successes. [You may find it useful to reference the z table.]

a. Construct the a 95% confidence interval for the population proportion of successes. (Round intermediate calculations to at least 4 decimal places. Round "z" value and final answers to 3 decimal places.)  


b. Construct the a 95% confidence interval for the population proportion of failures. (Round intermediate calculations to at least 4 decimal places. Round "z" value and final answers to 3 decimal places.)
Answer #1

Solution :

Given that,

a) Point estimate = sample proportion = \hat p = x / n = 104 / 160 = 0.65

1 - \hat p = 1 - 0.65 = 0.35

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha / 2 * ((\hat p * (1 - \hat p ))\sqrt / n)

= 1.96 (\sqrt((0.65 * 0.35) / 160)

= 0.074

A 95% confidence interval for population proportion p is ,

\hat p ± E  

= 0.65  ± 0.074

= ( 0.576, 0.724 )

b) x = 160 - 104 = 56

Point estimate = sample proportion = \hat p = x / n = 56 / 160 = 0.35

1 - \hat p = 1 - 0.35 = 0.65

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha / 2 * ((\hat p * (1 - \hat p ))\sqrt / n)

= 1.96 (\sqrt((0.35 * 0.65) / 160)

= 0.074

A 95% confidence interval for population proportion p is ,

\hat p ± E  

= 0.35  ± 0.074

= ( 0.276, 0.424 )